# Wireless Datasets for Aerial Networks

### v1

@@ V-A Description of Hardware and Software
Figure 1: Illustration of the AERPAW large multirotor-type UAV setup for the experiment, where the UAV carries a portable node [63].
The large multirotor-type UAV from AERPAW was used to collect I/Q samples during the experiment. As shown in Fig. 1, the UAV carries a portable payload that includes an NI USRP B205mini SDR. Python scripts run on the SDR and GPS module to collect I/Q samples at the desired center frequency, sampling rate, and interval, and to record the UAV’s location and position. The UAV is equipped with a dipole-type antenna (SA-1400-5900). Before the experiment, the UAV’s flight path, navigation speed, and position were pre-planned by placing waypoints on the map, enabling automatic control of the UAV and repeatable experiments. The collected I/Q dataset is stored in MATLAB (.mat) format and can be post-processed using MATLAB toolboxes. The GPS logs are also saved in a text file (.txt) for post-processing.
An AERPAW fixed radio node at the Lake Wheeler Road Field Labs (LWRFL) site (see Fig. 2(a) and Fig. 2(b)) is configured as an LTE Evolved Node B (eNB) to transmit the LTE downlink signal. The srsRAN open-source SDR software is used to realize the LTE eNB, where the transmitter antenna gain, center frequency, and number of resource blocks are configurable. A USRP B205mini SDR with a wideband antenna (RM-WB1) is installed at the fixed radio node.
@@ V-B Dataset Format
((a))
((b))
((c))
((d))
@@
The I/Q measurement campaign was conducted using an LTE BS and a UAV (see Fig. 2(c)), which followed a zigzag trajectory at five different fixed altitudes, as illustrated in Fig. 2(d). The altitudes ranged from $30$ m to $110$ m, with increments of $20$ m between each. The I/Q samples and the UAV’s GPS logs datasets are available on IEEE Dataport [24], providing a valuable resource for in-depth research involving raw I/Q data. The dataset is organized into five main folders corresponding to the I/Q measurements taken at each altitude. Each folder is named according to the respective altitude. For every altitude, the data is further divided into two subfolders: one for the I/Q samples and the other for the GPS logs. The raw data from both sources are converted into SigMF [64], accompanied by metadata that describes the recordings information. Scripts are provided to facilitate conversion into raw data formats compatible with MATLAB or Python (e.g., .mat or comma-separated values (CSV) files). For each altitude, the dataset is structured as follows:
  1. 1.
    I/Q Sample Data: These files are in SigMF format and are located in the IQ_samples folder. Each file contains I/Q samples collected over a duration of $20$ milliseconds, and the files are named according to the timestamp of the I/Q sample measurement start time. Consecutive files are separated by $100$ milliseconds, corresponding to the interval between measurements. With a $2$ MHz sampling rate, each SigMF file contains approximately $40k$ I/Q samples. Additionally, the measurement collection time in Unix timestamp is included within the SigMF data.
  2. 2.
    GPS Logs: The GPS log file in SigMF format is stored in the GPS_logs folder. When converted to CSV format, each row includes the latitude, longitude, altitude (in meters), and the Unix timestamp of the GPS measurement. The record interval is one second.
To facilitate the conversion of SigMF-formatted files into more widely used formats, the repository also provides the following Python scripts:
  • sigMF2mat_IQ.py: Converts I/Q sample data to MATLAB format. The output files are saved in the ./IQ_samples/matfile directory.
  • sigMF2csv_GPS.py: Converts GPS log file into CSV format. The output file is saved in the ./GPS_logs/csvfile directory.
((a))
((b))
((a))
((b))
((c))
((d))
((e))
((f))
Figure 4: Representative results from the Wireless I/Q dataset: (a) LTE resource grid, (b) estimated channel, (c) RSRP along UAV trajectory, (d) time vs. 3D distance (UAV altitude: $70\,\mathrm{m}$), (e) time vs. UAV speed (UAV altitude: $70\,\mathrm{m}$), and (f) 3D distance vs. RSRP with path loss fitting (UAV altitude: $70\,\mathrm{m}$).
Additionally, a README file is provided, offering an overview of dataset usage. The dataset is also equipped with a detailed description of the methodology, equipment, environment, and post-processing codes for air-to-ground (A2G) propagation modeling [65]. Post-processing scripts are available on CodeOcean [25] for extracting channel information, coherence bandwidth, RSRP, and RSRQ from the I/Q samples.
@@ V-C Representative Results
In this section, we present several representative results using the published post-processing scripts [25, 26]. Since the I/Q samples in this dataset are LTE waveforms and the measurement window is $20$ milliseconds, we can decode the LTE frame start time and extract a full $10$ millisecond LTE frame from each I/Q measurement MATLAB file. We can then plot the LTE resource grid and estimate the channel using the reference signals, as shown in Fig. 4(a) and Fig. 4(b), respectively. From each I/Q measurement file, we can also obtain the RSRP value, and by matching the Unix timestamp with the GPS logs, we can determine the corresponding 3D location of the UAV and plot the RSRP over the UAV trajectory, as shown in Fig. 4(c). Additionally, from the GPS logs, we can plot the 3D distance from the BS to the UAV over time and calculate the UAV’s speed by differentiating the position, as shown in Fig. 4(d) and Fig. 4(e), respectively. Finally, Fig. 4(f) accumulates the GPS logs and RSRP data along the Unix timestamp to plot the RSRP across the 3D distance, where the fitted free-space and two-ray path loss models incorporating 3D antenna patterns are aligned with the measured RSRP behavior.
@@ V-D Possible Uses of Dataset
V-D1 A2G Propagation Modeling
Artificial intelligence (AI)-based A2G propagation models can be developed using the provided I/Q datasets. AI-based training and testing models can incorporate multiple features, such as the 3D antenna pattern, UAV altitude and position, BS tower height, communication frequency and bandwidth, UAV speed. Note that our dataset [66] includes 3D radiation pattern measurements of both the transmit and receive antennas, obtained in an anechoic chamber.
V-D2 UAV Receiver Algorithm Design
The provided I/Q dataset can facilitate the development of practical time and frequency synchronization algorithms, cell search, channel estimation, and decoding techniques optimized for UAV communication systems.
V-D3 Spectrum Occupancy and Interpolation
For spectrum sharing and coexistence between terrestrial and aerial networks, UAV-based spectrum monitoring and interpolation techniques have been widely investigated. Our I/Q dataset includes a single zigzag trajectory at multiple altitudes, enabling altitude-dependent spectrum analysis and the study of spectrum interpolation techniques. In our preliminary works [27, 67], we propose spectrum interpolation approaches based on the 3D Kriging [27] and matrix completion [67] using the I/Q dataset.
V-D4 UAV Localization and Tracking
The detection, localization, and tracking of signal sources by UAVs are key techniques for ensuring privacy and enabling network coexistence. By classifying malicious UAVs or incumbent signal sources, UAV-based systems can enhance situational awareness and support secure and reliable spectrum operations. Our I/Q dataset provides received signal strength (RSS) measurements along the UAV trajectory across the experiment site, enabling the development and evaluation of source localization and tracking algorithms. In our preliminary work, we propose a UAV localization technique based on the two-ray path loss model, incorporating 3D antenna radiation patterns [28].
@@ VI-A Description of Hardware and Software
((a))
((b))
Figure 5: Measurement setup and procedure for spectrum data collection using the helikite-mounted portable node. (a) Experimental setup of the portable node on the tethered helikite. (b) Spectrum sweep procedure.
Fig. 5(a) illustrates the equipment configuration attached to the helikite. On the left, the helikite adapter is shown, which houses the battery and the GPS logger. On the right, the adapter accommodates the portable node. The battery supplies power to both the portable node and the GPS logger. GPS logger is a very simple AERPAW sample application that only logs the location of the vehicle (based on information from the vehicle’s autopilot) but does not control the vehicle at all. The purpose of this sample application is to allow the monitoring of the vehicle while the vehicle is controlled externally. This allows for drones and rover to be operated under manual control from the pilot, while still logging their position. The GPS logger is connected via a USB cable to a companion computer (Intel NUC) housed within the portable node. The Intel NUC manages two B205mini-i SDRs; one is equipped with a 3.5 GHz front end while the other operates without a specific front end.
The portable node is designed to operate on a 19 V supply from the battery, which powers not only the companion computer but also the active components of the front end, such as the Low Noise Amplifier (LNA). Both USRPs are connected to identical but separate receive antennas and are configured only to receive signals; there is no transmission functionality in this setup. The measured data for each sweep requires approximately 15 seconds to be stored. The spectrum sweep procedure used in the experiments is depicted in Fig. 5(b), where the center frequency shift and sampling rate are 25.68 MHz and 30.72 MHz, respectively. This dataset presents results exclusively for the entire bandwidth, i.e., from 87 MHz to 6 GHz.
@@ VI-B Dataset Format
The dataset available on the Dryad Digital Repository is comprehensively structured to support extensive research into spectrum analysis and wireless communications. It is organized into several key components aimed at providing a robust set of tools for data analysis and application.
The dataset comprises three primary types of files stored in designated folders:
  1. 1.
    Power Spectrum Raw Data: These are SigMF formatted files located in the pow_spec folder. Each MATLAB file contains frequency data (freqs in MHz) and power measurements (powers in dBm), with filenames that correspond to timestamps indicative of the data collection time.
  2. 2.
    GPS Logs: SigMF formatted GPS logs are stored in the GPS_logs folder. These logs are convertible to CSV format using provided Python scripts. The converted CSV files include four columns: longitude, latitude, altitude (in meters), and Unix epoch timestamp.
Additionally, the repository provides Python scripts to facilitate the conversion of these SigMF formatted files into more widely used formats:
  • sigMF2mat_PW.py: Converts power spectrum data to MATLAB format. The output files are saved in the ./pow_spec/matfile directory.
  • sigMF2csv_GPS.py: Converts GPS logs into CSV format. The output files are stored in the ./GPS_logs/csvfile directory.
To use these scripts, users can run sigMF2mat_PW.py to process the power spectrum data, and sigMF2csv_GPS.py to extract the GPS logs. Each script ensures that data is accessible and easy to manipulate, enhancing the usability of the dataset for various research purposes.
This dataset is equipped with detailed metadata, which includes the methodology, equipment used, settings, and examples of post-processing. This comprehensive metadata ensures that users can reliably replicate measurements and fully understand the legal and ethical considerations for using the data. Additionally, a README file is included, providing a comprehensive overview of the dataset’s scope, usage instructions, and curator contact information. This essential support is invaluable for effectively interpreting and applying the data to specific research needs.
@@ VI-C Representative Results
((a))
((b))
Figure 7: Helikite location for spectrum measurements in (a) Packapalooza 2024 and (b) Lake Wheeler 2024.
VI-C1 Helikite Trajectory Analysis
In urban settings during the Packapalooza event, the helikite’s trajectory, as depicted in 7(a), showcases its path above a densely populated area with significant deviations caused by complex wind interactions with urban structures. This erratic movement, indicated by a red trace, potentially affects spectrum measurements due to variable altitudes and obstructions. Conversely, in the rural landscape near Lake Wheeler as shown in Fig. 7(b), the helikite exhibits a more stable and elongated flight path across open fields, suggesting more consistent data collection due to fewer obstructions and a steadier altitude control.
((a))
((b))
Figure 8: Altitude dependent power for DL 13 in (a) Packapalooza 2024 and (b) Lake Wheeler 2024.
@@ VI-C1 Helikite Trajectory Analysis
VI-C2 Spectrum Analysis
In Fig. 8, the 3D plots for Band 13 downlink (746 - 756 MHz) reveal distinct variations in signal power across different environments. In urban settings, Fig. 8(a), power levels fluctuate significantly with altitude due to multipath effects and obstructions, showing a trend toward stabilization as altitude increases. In contrast, rural settings in Fig. 8(b) display a more uniform increase in power levels at higher altitudes, indicating clearer signal paths and fewer obstructions. Fig. 9 illustrates the mean received power as a function of frequency across altitudes for the Packapalooza 2024 dataset, with several active United States radio bands also indicated. The results show that the mean received power is significantly higher below 1 GHz, where many LTE and NR network bands are located, compared to the remainder of the sub-6 GHz spectrum.
@@ VI-D1 Spectrum Allocation Analysis
The dataset enables spatial and altitudinal characterization of spectrum utilization across urban and rural environments. For example, comparisons of received power in the Citizens Broadband Radio Service (CBRS) and Television White Spaces (TVWS) bands can reveal underutilized areas or high-demand regions. In urban settings, it allows researchers to measure outdoor signal levels in the 6 GHz unlicensed band to assess potential interference from indoor Wi-Fi 6E deployments. The availability of time-stamped data also permits the exploration of temporal usage trends, such as peak usage periods or band-specific congestion.
VI-D2 Calibration of Analytical Models
@@
The dataset supports calibration of analytical and simulation models, including stochastic geometry and empirical path loss frameworks. The relationship between received power and altitude can be used to validate altitude-aware propagation assumptions. Differences observed between rural and urban measurements can help refine clutter loss models. In addition, LoS probability models can be empirically evaluated using elevation-dependent signal trends.
VI-D3 Propagation Model Tuning
@@
Researchers can use the dataset to develop or refine radio propagation models. The location-tagged power measurements support construction of empirical path loss curves for a range of frequency bands. Comparing measurements taken in urban versus rural environments helps to characterize the impact of buildings, vegetation, and other obstructions. Moreover, differences between low-frequency bands like FM and higher bands like 3.5 GHz can be used to study frequency-dependent attenuation.
VI-D4 Machine Learning for Signal Estimation
@@
The dataset is suitable for developing machine learning models that estimate received power from spatial and environmental features. Inputs such as latitude, longitude, altitude, and frequency can be used to train regressors for power prediction. The labeled nature of rural and urban environments supports classification tasks, such as identifying the type of environment based on observed signal levels. The data can also help delineate signal boundaries or approximate coverage maps through supervised learning.
VI-D5 Anomaly and Interference Detection
@@
The received power measurements allow for basic anomaly detection techniques. Sudden spikes or dips in power levels may indicate unauthorized transmissions or interference events. Statistical properties such as skewness or variance can be used to detect deviations from normal signal patterns. When the data is tracked over time, researchers can analyze signal disruptions or temporal anomalies in specific frequency bands.
VI-D6 Interpolation and Coverage Mapping
@@
Despite the absence of I/Q or PSD data, the dataset is well-suited for generating radio environment maps. The geolocation and altitude information associated with each measurement can be used for spatial interpolation techniques such as Kriging or inverse distance weighting. By interpolating the power values, researchers can generate two-dimensional (2D) or 3D signal coverage maps for individual bands. The multi-band nature of the dataset further allows for frequency-aware coverage visualizations across the monitored spectrum.
@@
Figure 10: Ericsson 5G-NSA AERPAW Infrastructure (top), and the SW workflow with 5G SPN on the 5G-NSA network (bottom).

### v2

@@ V-A Description of Hardware and Software
Figure 1: Illustration of the AERPAW large multirotor-type UAV setup for the experiment, where the UAV carries a portable node.
The large multirotor-type UAV from AERPAW was used to collect I/Q samples during the experiment. As shown in Fig. 1, the UAV carries a portable payload that includes an NI USRP B205mini SDR. Python scripts run on the SDR and GPS module to collect I/Q samples at the desired center frequency, sampling rate, and interval, and to record the UAV’s location and position. The UAV is equipped with a dipole-type antenna (SA-1400-5900). Before the experiment, the UAV’s flight path, navigation speed, and position were pre-planned by placing waypoints on the map, enabling automatic control of the UAV and repeatable experiments. The released dataset includes the recorded I/Q samples and the corresponding GPS logs. The raw data are stored in MATLAB (.mat) format, while the GPS information is provided as text files (.txt) to facilitate post-processing.
An AERPAW fixed radio node at the Lake Wheeler Road Field Labs (LWRFL) site (see Fig. 2a and Fig. 2b) is configured as an LTE Evolved Node B (eNB) to transmit the LTE downlink signal. The srsRAN open-source SDR software is used to realize the LTE eNB, where the transmitter antenna gain, center frequency, and number of resource blocks are configurable. A USRP B205mini SDR with a wideband antenna (RM-WB1) is installed at the fixed radio node.
@@ V-B Dataset Format
(a) Campaign Location
(b) Base Station
(c) AERPAW UAV
(d) Trajectory
@@
The I/Q measurement campaign was conducted using an LTE base station and a UAV platform (see Fig. 2c), where the UAV followed a zigzag trajectory at five fixed altitudes ranging from $30$ m to $110$ m in $20$ m increments, as illustrated in Fig. 2d. The resulting wireless I/Q samples and synchronized UAV GPS logs are publicly available on IEEE Dataport [52], providing a high-fidelity resource for air-to-ground (A2G) propagation analysis, channel estimation, and signal-level modeling.
All measurements are released in SigMF format [85], including standardized metadata that captures recording parameters, timing information, and measurement context. The I/Q recordings consist of consecutive $20$ ms snapshots acquired at a $2$ MHz sampling rate, while GPS logs provide time-stamped latitude, longitude, and altitude information at one-second intervals. These data enable precise alignment between signal-level observations and UAV mobility.
For reproducibility and ease of reuse, the dataset is accompanied by documentation and post-processing utilities that support conversion to commonly used analysis formats such as MATLAB and CSV. Detailed directory hierarchy, file naming conventions, and conversion script descriptions are provided in Appendix A.
(a) LTE Resource Grid
(b) Estimated time and frequency selective channel.
(c) RSRP at five altitudes.
(d) Time vs 3D Distance
(e) Time vs UAV Speed
(f) Measured and modeled RSRP vs. 3D distance.
Figure 3: Representative results from the Wireless I/Q dataset: (a) LTE resource grid, (b) estimated channel, (c) RSRP along UAV trajectory, (d) time vs. 3D distance (UAV altitude: $70\,\mathrm{m}$), (e) time vs. UAV speed (UAV altitude: $70\,\mathrm{m}$), and (f) 3D distance vs. RSRP with path loss fitting (UAV altitude: $70\,\mathrm{m}$). The figure shows that RSRP decreases as cumulative distance increases, while the speed and distance traces confirm measurements were taken under well-controlled mobility, supporting statistically consistent path-loss and fading analysis.
@@ V-C Representative Results
In this section, we present several representative results using the published post-processing scripts [54, 60]. Since the I/Q samples in this dataset are LTE waveforms and the measurement window is $20$ milliseconds, we can decode the LTE frame start time and extract a full $10$ millisecond LTE frame from each I/Q measurement MATLAB file. We can then plot the LTE resource grid and estimate the channel using the reference signals, as shown in Fig. 3a and Fig. 3b, respectively. From each I/Q measurement file, we can also obtain the RSRP value, and by matching the Unix timestamp with the GPS logs, we can determine the corresponding 3D location of the UAV and plot the RSRP over the UAV trajectory, as shown in Fig. 3c. Additionally, from the GPS logs, we can plot the 3D distance from the BS to the UAV over time and calculate the UAV’s speed by differentiating the position, as shown in Fig. 3d and Fig. 3e, respectively. Finally, Fig. 3f accumulates the GPS logs and RSRP data along the Unix timestamp to plot the RSRP across the 3D distance, where the fitted free-space and two-ray path loss models incorporating 3D antenna patterns are aligned with the measured RSRP behavior.
@@ V-D Possible Uses of Dataset
A2G Propagation Modeling: Artificial intelligence (AI)-based A2G propagation models can be developed using the provided I/Q datasets. AI-based training and testing models can incorporate multiple features, such as the 3D antenna pattern, UAV altitude and position, BS tower height, communication frequency and bandwidth, UAV speed. Note that our dataset [56] includes 3D radiation pattern measurements of both the transmit and receive antennas, obtained in an anechoic chamber.
UAV Receiver Algorithm Design: The provided I/Q dataset can facilitate the development of practical time and frequency synchronization algorithms, cell search, channel estimation, and decoding techniques optimized for UAV communication systems.
Spectrum Occupancy and Interpolation: For spectrum sharing and coexistence between terrestrial and aerial networks, UAV-based spectrum monitoring and interpolation techniques have been widely investigated. Our I/Q dataset includes a single zigzag trajectory at multiple altitudes, enabling altitude-dependent spectrum analysis and the study of spectrum interpolation techniques. In our preliminary works [55, 72], we propose spectrum interpolation approaches based on the 3D Kriging [55] and matrix completion [72] using the I/Q dataset.
UAV Localization and Tracking: The detection, localization, and tracking of signal sources by UAVs are key techniques for ensuring privacy and enabling network coexistence. By classifying malicious UAVs or incumbent signal sources, UAV-based systems can enhance situational awareness and support secure and reliable spectrum operations. Our I/Q dataset provides received signal strength (RSS) measurements along the UAV trajectory across the experiment site, enabling the development and evaluation of source localization and tracking algorithms. In our preliminary work, we propose a UAV localization technique based on the two-ray path loss model, incorporating 3D antenna radiation patterns [51].
@@ VI-A Description of Hardware and Software
(a) Setup of the portable node on the helikite.
(b) Spectrum sweep procedure.
Figure 4: Measurement setup and procedure for spectrum data collection using the helikite-mounted portable node. (a) Experimental setup of the portable node on the tethered helikite. (b) Spectrum sweep procedure. The procedure indicates that spectrum measurements are collected using repeatable sweeps under fixed hardware settings, making the resulting power statistics suitable for comparative altitude- and time-based analysis.
Fig. 4a illustrates the helikite-mounted measurement configuration, including the adapter housing the battery, GPS logger, and portable node. The GPS logger records the platform location using information from the vehicle autopilot while the platform is externally controlled. The GPS logger is connected via a USB cable to a companion computer (Intel NUC) housed within the portable node. The Intel NUC manages two B205mini-i SDRs; one is equipped with a 3.5 GHz front end while the other operates without a specific front end.
The portable node is designed to operate on a 19 V supply from the battery, which powers not only the companion computer but also the active components of the front end, such as the Low Noise Amplifier (LNA). Both USRPs are connected to identical but separate receive antennas and are configured only to receive signals; there is no transmission functionality in this setup. The measured data for each sweep requires approximately 15 seconds to be stored. The spectrum sweep procedure used in the experiments is depicted in Fig. 4b, where the center frequency shift and sampling rate are 25.68 MHz and 30.72 MHz, respectively.
@@ VI-B Dataset Format
The wireless spectrum measurement dataset is publicly available through the Dryad Digital Repository and is designed to support wideband spectrum analysis, propagation studies, and data-driven wireless research. The dataset consists of time-stamped power spectral measurements collected during aerial experiments, together with synchronized UAV location information.
All spectrum and positioning data are released in SigMF format, accompanied by standardized metadata that document the measurement methodology, equipment configuration, frequency settings, and acquisition context. Power spectrum records provide frequency-domain power measurements expressed in dBm over the monitored bandwidth, while GPS logs include time-aligned latitude, longitude, altitude, and Unix timestamp information. This structure enables precise association between spectrum observations and UAV mobility.
To facilitate reproducible analysis and interoperability with commonly used signal processing tools, the dataset includes supporting documentation and utilities for conversion into MATLAB and CSV formats. Detailed directory hierarchy, file naming conventions, and conversion script descriptions are provided in Appendix A.
@@ VI-C Representative Results
(a) Packapalooza 2024.
(b) Lake Wheeler 2024.
Figure 5: Helikite location for spectrum measurements in (a) Packapalooza 2024 and (b) Lake Wheeler 2024. The two locations illustrate that the dataset captures environment-dependent spectrum statistics, enabling practical comparison between less congested and more interference-prone RF environments.
(a) Packapalooza 2024.
(b) Lake Wheeler 2024.
Figure 6: Altitude dependent power in the 746756746-756 spectrum (LTE DL Band-13) in (a) Packapalooza 2024 and (b) Lake Wheeler 2024. The plot shows systematic power fluctuations with altitude, indicating that elevation is a statistically significant factor in received spectrum power for aerial platforms.
@@
Helikite Trajectory Analysis: In urban settings during the Packapalooza event, the helikite’s trajectory, as depicted in Fig. 5a, showcases its path above a densely populated area with significant deviations caused by complex wind interactions with urban structures. This erratic movement, indicated by a red trace, potentially affects spectrum measurements due to variable altitudes and obstructions. Conversely, in the rural landscape near Lake Wheeler as shown in Fig. 5b, the helikite exhibits a more stable and elongated flight path across open fields, suggesting more consistent data collection due to fewer obstructions and a steadier altitude control.
Spectrum Analysis: In Fig. 6, the 3D plots for Band 13 downlink (746 - 756 MHz) reveal distinct variations in signal power across different environments. In urban settings, Fig. 6a, power levels fluctuate significantly with altitude due to multipath effects and obstructions, showing a trend toward stabilization as altitude increases. In contrast, rural settings in Fig. 6b display a more uniform increase in power levels at higher altitudes, indicating clearer signal paths and fewer obstructions. Fig. 7 illustrates the mean received power as a function of frequency across altitudes for the Packapalooza 2024 dataset, with several active United States radio bands also indicated. The results show that the mean received power is significantly higher below 1 GHz, where many LTE and NR network bands are located, compared to the remainder of the sub-6 GHz spectrum.
@@ VI-D Possible Uses of Dataset
Spectrum Allocation Analysis: The dataset enables spatial and altitudinal characterization of spectrum utilization across urban and rural environments. For example, comparisons of received power in the Citizens Broadband Radio Service (CBRS) and Television White Spaces (TVWS) bands can reveal underutilized areas or high-demand regions. In urban settings, it allows researchers to measure outdoor signal levels in the 6 GHz unlicensed band to assess potential interference from indoor Wi-Fi 6E deployments. The availability of time-stamped data also permits the exploration of temporal usage trends, such as peak usage periods or band-specific congestion.
@@
Calibration of Analytical Models: The dataset supports calibration of analytical and simulation models, including stochastic geometry and empirical path loss frameworks. The relationship between received power and altitude can be used to validate altitude-aware propagation assumptions. Differences observed between rural and urban measurements can help refine clutter loss models. In addition, LoS probability models can be empirically evaluated using elevation-dependent signal trends.
@@
Propagation Model Tuning: Researchers can use the dataset to develop or refine radio propagation models. The location-tagged power measurements support construction of empirical path loss curves for a range of frequency bands. Comparing measurements taken in urban versus rural environments helps to characterize the impact of buildings, vegetation, and other obstructions. Moreover, differences between low-frequency bands like FM and higher bands like 3.5 GHz can be used to study frequency-dependent attenuation.
@@
Machine Learning for Signal Estimation: The dataset is suitable for developing machine learning models that estimate received power from spatial and environmental features. Inputs such as latitude, longitude, altitude, and frequency can be used to train regressors for power prediction. The labeled nature of rural and urban environments supports classification tasks, such as identifying the type of environment based on observed signal levels. The data can also help delineate signal boundaries or approximate coverage maps through supervised learning.
@@
Anomaly and Interference Detection: The received power measurements allow for basic anomaly detection techniques. Sudden spikes or dips in power levels may indicate unauthorized transmissions or interference events. Statistical properties such as skewness or variance can be used to detect deviations from normal signal patterns. When the data is tracked over time, researchers can analyze signal disruptions or temporal anomalies in specific frequency bands.
@@
Interpolation and Coverage Mapping: Despite the absence of I/Q or PSD data, the dataset is well-suited for generating radio environment maps. The geolocation and altitude information associated with each measurement can be used for spatial interpolation techniques such as Kriging or inverse distance weighting. By interpolating the power values, researchers can generate two-dimensional (2D) or 3D signal coverage maps for individual bands. The multi-band nature of the dataset further allows for frequency-aware coverage visualizations across the monitored spectrum.
@@
Figure 8: 5G-NSA AERPAW Infrastructure using Ericsson network.